calculate the pH in a solution prepared by dissolving 0.10 mol of NH4CH3COO in 0.500Lof 0.40 M HC2H3O2. Ka=1.8*10^-5

please show work
Answer #1

pH = pKa + log(salt/acid)

     = 4.745 + log( 0.1/0.4)

     = 4.745 - 0.6

     = 4.145